# Visibility-Aware Language Aggregation for Open-Vocabulary Segmentation in 3D Gaussian Splatting

### v1

@@ Appendix B Evaluation Protocols
Datasets We evaluate our method on two datasets: LERF-OVS [27] and ScanNet [5]. LERF-OVS consists of four scenes (teatime, waldo_kitchen, figurines, ramen), each annotated with pixel-wise semantic masks and paired with short text queries. On this dataset, we assess open-vocabulary object selection in both 2D and 3D. To further evaluate 3D semantic segmentation, we adopt a Gaussian-based evaluation protocol on ScanNet, a large-scale RGB-D dataset for indoor scene understanding. Each ScanNet sequence is reconstructed into a textured 3D mesh with globally aligned camera poses and semantic annotations. We select eight representative scenes covering diverse indoor environments, including living rooms, bathrooms, kitchens, bedrooms, and meeting rooms.
@@
2D and 3D Evaluation on the LERF-OVS Dataset. For the 2D evaluation, we follow the protocol of LERF [14]: 512-dimensional feature maps are rendered, and a relevancy map with respect to the CLIP-embedded text query is computed. The relevancy map is then thresholded at 0.5 to obtain the predicted binary mask. For the 3D evaluation, we adopt the protocol of OpenGaussian [37], where the relevancy score between each 3D Gaussian’s language embedding and the text query embedding is computed and thresholded at 0.6. The alpha values of the selected Gaussians are subsequently projected onto the image plane to generate the predicted mask. In both cases, the predicted masks are compared against the GT annotations of the LERF-OVS dataset.
@@
3D Semantic Segmentation on the ScanNet-v2 Dataset. Previous protocols [37] freeze the input point cloud during evaluation, which reduces rendering fidelity. Inspired by Dr.Splat [12], we instead propagate ground-truth (GT) labels from the annotated point cloud to the Gaussians, thereby obtaining pseudo-GT labels at each Gaussian’s 3D mean. Following OpenGaussian [37], we evaluate on subsets of 19, 15, and 10 of the 40 most common classes. For each class, we encode the text label using CLIP [29] to obtain a 512-dimensional embedding, and compute its cosine similarity with the registered language features of each Gaussian. Each Gaussian is then assigned to the class with the highest similarity score. Performance is measured in terms of mIoU and mAcc against the pseudo-GT Gaussian point cloud.
Pseudo-Gaussian Labeling. Previous works on 3D semantic segmentation typically freeze the input point cloud (derived from ground-truth annotations) during 3D Gaussian Splatting training to cope with the absence of GT labels as the point clouds evolve. However, this strategy degrades the 2D rendering quality of 3DGS. Inspired by Kim et al. [12], we instead first train 3D Gaussian Splatting without restrictions and subsequently propagate labels from the ground-truth point cloud to the optimized Gaussians, explicitly accounting for their scales and rotations.
@@
Given optimized Gaussians $\Theta=\{\theta_{i}\}_{i=1}^{N}$ with center $\mu_{i}$, scale $s_{i}=(s_{ix},s_{iy},s_{iz})$, rotation $R_{i}$ (hence $\Sigma_{i}=R_{i}\,\mathrm{diag}(s_{i}^{2})\,R_{i}^{\top}$), and opacity $\alpha_{i}$, and a labeled point cloud $\{(p_{k},s_{p_{k}})\}_{k=1}^{Q}$, we assign a semantic label to each Gaussian by respecting the true 3DGS geometry and the compositing kernel. In contrast to prior protocols, which (i) maximize the sum of Mahalanobis distances over class points to assign a single label, and (ii) require dense all-pairs computations, our approach assigns semantic labels by respecting the true 3DGS geometry and properties. Specifically, we evaluate the density contribution of a point $p$ to Gaussian $\mu_{i}$:
@@
$\gamma_{i}=\alpha_{i}\,s_{ix}s_{iy}s_{iz},\quad\quad w_{i}(p)\leftarrow\gamma_{i}\,w_{i}(p).$ (19)
@@ Appendix C Robustness Evaluation with Perturbed Masks
When boundary damage becomes severe, however, the picture changes. VALA (w/o G) overtakes the full VALA model in accuracy (e.g., at $r=30$, achieving 9.95/15.25 vs. 6.75/1.69 in mIoU/mAcc) and consistently yields the lowest Disp across all radii. This suggests that the fixed gating threshold becomes overly conservative under extreme corruption, discarding too many observations and leaving insufficient evidence for many Gaussians. In contrast, the cosine-median aggregator alone remains robust, preserving both accuracy and consistency in this challenging setting. Overall, these results highlight a clear regime split: visibility-aware gating combined with a cosine median provides the strongest accuracy and consistency under realistic (mild–moderate) noise, whereas under extreme boundary corruption, robust aggregation is the key factor as overly strict gating thresholds reduce coverage and hurt performance.
@@ Appendix D Additional Results
More Qualitative Results on the ScanNet Dataset. We provide additional qualitative results on three bedroom scenes with varying levels of complexity and clutter. Across all scenes, competing methods struggle to correctly recognize the bed (highlighted in orange): the occluded portions near the wall are consistently misclassified as adjacent categories such as wall or floor. This issue persists in the third scene, where the bed is fragmented into multiple categories. In contrast, our method preserves the bed as a coherent instance, owing to the proposed gating module that explicitly handles low-visibility Gaussians.
Experiments on the Waymo Open Dataset. To further validate our algorithm’s generalization capability in real-world outdoor environments, we conduct experiments on the Waymo Open Dataset [32]. This dataset is a large-scale, high-quality autonomous driving benchmark that provides synchronized LiDAR and multi-camera data collected across diverse urban and suburban geographies, along with comprehensive 2D/3D annotations and tracking identifiers. For evaluation, we select a sequence captured in a residential neighborhood, which contains rich semantic elements such as vehicles, vegetation, street infrastructure, and buildings. We focus on five of the most common outdoor categories—tree, trash bin, car, streetlight, and house—as well as one tail category, stair. The qualitative results in Figure 8 demonstrate that our method achieves precise open-vocabulary 3D semantic segmentation on outdoor data. Both small-scale objects (e.g., trash bins and streetlights) and large-scale objects (e.g., trees, cars, and houses) are not only correctly retrieved but also segmented with sharp boundaries, reflecting the accurate registration of language features on the 3D Gaussian Splatting representation. Notably, our method remains robust under occlusion—for example, correctly delineating trees behind metallic structures or houses partially obscured by vegetation—owing to the proposed visibility-aware gating module. competing baselines frequently misclassify small, occluded, or distant objects (e.g., trash bins blend into the background; tree foliage gets mislabeled as building or sky). In contrast, our model consistently produces sharp, semantically coherent outputs: street lights, trash bins, and other fine objects are accurately recognized, while larger objects such as cars and houses remain seamlessly segmented.
These findings emphasize the robustness and versatility of our method when transferred from indoor (ScanNet) to challenging outdoor driving scenarios, underscoring its strong potential for real-world autonomous driving applications. A supplementary video is included to further demonstrate the effectiveness and the multi-view consistency of our method. More details can be found in the attached material.
@@ Appendix E Limitations
While our approach demonstrates strong performance across multiple tasks, including 2D and 3D object selection as well as 3D semantic segmentation, and exhibits notable generalization to cross-domain settings such as outdoor datasets, certain limitations remain. To assess robustness against noisy SAM masks, we conducted stress tests with multi-scale morphological perturbations. The results show that our visibility-aware gating achieves superior mIoU and mAcc under moderate noise, while the proposed cosine median maintains low dispersion even under severe corruption, indicating the effectiveness of our robust feature aggregator. However, our current framework relies on a fixed threshold to prune Gaussians, which may become over-conservative under extreme noise, leading to degraded multi-view consistency. Moreover, our method is designed for static scenes and does not extend naturally to dynamic environments. Future work will therefore focus on developing adaptive, scene-aware thresholds and extending our framework to handle dynamic scenes.

### v2

@@ Appendix B Evaluation Protocols
Datasets We evaluate our method on two datasets: LERF-OVS [28] and ScanNet [5]. LERF-OVS consists of four scenes (teatime, waldo_kitchen, figurines, ramen), each annotated with pixel-wise semantic masks and paired with short text queries. In this dataset, we evaluate open-vocabulary object selection in both 2D and 3D settings. To further evaluate 3D semantic segmentation, we adopt a Gaussian-based evaluation protocol on ScanNet, a large-scale RGB-D dataset for indoor scene understanding. Each ScanNet sequence is reconstructed into a textured 3D mesh with globally aligned camera poses and semantic annotations. We select eight representative scenes covering diverse indoor environments, including living rooms, bathrooms, kitchens, bedrooms, and meeting rooms.
@@
2D and 3D Evaluation on the LERF-OVS Dataset. For the 2D evaluation, we follow the protocol of LERF [15]: 512-dimensional feature maps are rendered, and a relevancy map is computed with respect to the CLIP-embedded text query. The relevancy map is then thresholded at 0.5 to obtain the predicted binary mask. For the 3D evaluation, we adopt the protocol of OpenGaussian [38], where the relevancy score between each 3D Gaussian’s language embedding and the text query embedding is computed and thresholded at 0.6. The alpha values of the selected Gaussians are subsequently projected onto the image plane to generate the predicted mask. In both cases, the predicted masks are compared against the GT annotations of the LERF-OVS dataset.
@@
3D Semantic Segmentation on the ScanNet-v2 Dataset. Previous works on 3D semantic segmentation [38, 18] typically freeze the input point cloud (derived from ground-truth annotations) during 3D Gaussian Splatting training to cope with the absence of GT labels as the point clouds evolve. However, this strategy degrades the 2D rendering quality of 3DGS. We instead propagate ground-truth (GT) labels from the annotated point cloud to the Gaussians, thereby obtaining pseudo-GT labels at each Gaussian’s 3D mean. Following OpenGaussian [38], we evaluate on subsets of 19, 15, and 10 of the 40 most common classes. For each class, we encode the text label using CLIP [30] to obtain a 512-dimensional embedding, and compute its cosine similarity with the registered language features of each Gaussian. Each Gaussian is then assigned to the class with the highest similarity score. Performance is measured in terms of mIoU and mAcc against the pseudo-GT Gaussian point cloud.
@@
Pseudo-Gaussian Labeling. Given optimized Gaussians $\Theta=\{\theta_{i}\}_{i=1}^{N}$ with center $\mu_{i}$, scale $s_{i}=(s_{ix},s_{iy},s_{iz})$, rotation $R_{i}$ (hence $\Sigma_{i}=R_{i}\,\mathrm{diag}(s_{i}^{2})\,R_{i}^{\top}$), and opacity $\alpha_{i}$, and a labeled point cloud $\{(p_{k},s_{p_{k}})\}_{k=1}^{Q}$, we assign a semantic label to each Gaussian by respecting the true 3DGS geometry and the compositing kernel. In contrast to prior protocols, which (i) maximize the sum of Mahalanobis distances over class points to assign a single label, and (ii) require dense all-pairs computations, our approach assigns semantic labels by respecting the true 3DGS geometry and properties. Specifically, we evaluate the density contribution of a point $p$ to the Gaussian $\mu_{i}$:
@@
$\gamma_{i}=\alpha_{i}\,s_{ix}s_{iy}s_{iz},\qquad\qquad w_{i}(p)\leftarrow\gamma_{i}\,w_{i}(p).$ (19)
@@ Appendix C Robustness Evaluation with Perturbed Masks
When boundary damage becomes severe, however, the picture changes. VALA (w/o G) overtakes the full VALA model in accuracy (e.g., at $r=30$, achieving 9.95/15.25 vs. 6.75/1.69 in mIoU/mAcc) and consistently yields the lowest Disp across all radii. This suggests that the fixed gating threshold becomes overly conservative under extreme corruption, discarding too many observations and leaving insufficient evidence for many Gaussians. In contrast, the cosine-median aggregator alone remains robust, preserving both accuracy and consistency in this challenging setting. Overall, these results highlight a clear regime split: visibility-aware gating combined with a cosine median provides the strongest accuracy and consistency under realistic (mild to moderate) noise. However, under extreme boundary corruption, robust aggregation is the key factor, as overly strict gating thresholds reduce coverage and performance.
@@ Appendix D Additional Results
More Qualitative Results on the ScanNet Dataset. We provide additional qualitative results on three bedrooms with varying levels of complexity and clutter. Across all scenes, competing methods struggle to correctly recognize the bed (highlighted in orange); the occluded portions near the wall are consistently misclassified as adjacent categories, such as the wall or floor. This issue persists in the third scene, where the bed is fragmented into multiple categories. In contrast, our method preserves the bed as a coherent instance, owing to the proposed gating module that explicitly handles low-visibility Gaussians.
Experiments on the Waymo Open Dataset. To further validate our algorithm’s generalization capability in real-world outdoor environments, we conduct experiments on the Waymo Open Dataset [33]. This dataset is a large-scale, high-quality autonomous driving benchmark that provides synchronized LiDAR and multi-camera data collected across diverse urban and suburban geographies, along with comprehensive 2D/3D annotations and tracking identifiers. For evaluation, we select a sequence captured in a residential neighborhood that contains rich semantic elements, such as vehicles, vegetation, street infrastructure, and buildings. We focus on five of the most common outdoor categories, e.g. tree, trash bin, car, streetlight, and house, as well as one tail category, stair. The qualitative results in Figure 8 demonstrate that our method achieves precise open-vocabulary 3D semantic segmentation on outdoor data. Both small-scale objects (e.g., trash bins and streetlights) and large-scale objects (e.g., trees, cars, and houses) are not only correctly retrieved but also segmented with sharp boundaries, reflecting the accurate registration of language features on the 3D Gaussian Splatting representation. Notably, our method remains robust under occlusion owing to the proposed visibility-aware gating module. For example, correctly delineating trees behind metallic structures or houses partially obscured by vegetation.
These findings emphasize the robustness and versatility of our method when transferred from indoor (ScanNet) to challenging outdoor driving scenarios, underscoring its strong potential for real-world autonomous driving applications. A supplementary video is included to further demonstrate the effectiveness and the multi-view consistency of our method.
@@ Appendix E Limitations
While our approach demonstrates strong performance across multiple tasks, including 2D and 3D object selection as well as 3D semantic segmentation, and exhibits notable generalization to cross-domain settings such as outdoor datasets, certain limitations remain. To assess robustness against noisy SAM masks, we conducted stress tests with multi-scale morphological perturbations. The results show that our visibility-aware gating achieves superior mIoU and mAcc under moderate noise, while the proposed cosine median maintains low dispersion even under severe corruption, indicating the effectiveness of our robust feature aggregator. However, our current framework relies on a fixed threshold to prune Gaussians, which can become overly conservative under extreme noise, resulting in degraded multi-view consistency. Moreover, our method is specifically designed for static scenes and does not naturally extend to dynamic environments. Future work will therefore focus on developing adaptive, scene-aware thresholds and extending our framework to handle dynamic scenes.